### v1

# A Simple and Effective $L_{2}$ Norm-Based Strategy for KV Cache Compression

Abstract
The deployment of large language models (LLMs) is often hindered by the extensive memory requirements of the Key-Value (KV) cache, especially as context lengths increase. Existing approaches to reduce the KV cache size involve either fine-tuning the model to learn a compression strategy or leveraging attention scores to reduce the sequence length. We analyse the attention distributions in decoder-only Transformers-based models and observe that attention allocation patterns stay consistent across most layers. Surprisingly, we find a clear correlation between the $L_{2}$ norm and the attention scores over cached KV pairs, where a low $L_{2}$ norm of a key embedding usually leads to a high attention score during decoding. This finding indicates that the influence of a KV pair is potentially determined by the key embedding itself before being queried. Based on this observation, we compress the KV cache based on the $L_{2}$ norm of key embeddings. Our experimental results show that this simple strategy can reduce the KV cache size by 50% on language modelling and needle-in-a-haystack tasks and 90% on passkey retrieval tasks without losing accuracy.
A Simple and Effective $L_{2}$ Norm-Based Strategy
for KV Cache Compression
Alessio Devoto [^1] Equal contribution. Yu Zhao [^1] Equal contribution. Simone Scardapane Pasquale Minervini Sapienza University of Rome The University of Edinburgh {alessio.devoto, simone.scardapane}@uniroma1.it {yu.zhao, p.minervini}@ed.ac.uk
Figure 1: ALr, as defined in Eq. 3, for each head and layer in Llama2-7b. A lower value means a higher correlation between $L_{2}$ norm and attention score.
1 Introduction
Handling long contexts is desirable for large language models (LLMs), as it allows them to perform tasks that require understanding long-term dependencies Liu et al. (2024); Fu et al. (2024); Chen et al. (2023); Staniszewski et al. (2023); Tworkowski et al. (2024). A key component for modelling long context is the KV cache, which stores the keys and values of past tokens in memory to avoid recomputing them during generation. However, processing long-context inputs often results in a high decoding latency since it requires repeatedly reading a potentially large KV cache from high-bandwidth memory (HBM) to the streaming multiprocessor (SM) during decoding (Fu, 2024). Consequently, the practical deployment of LLMs is frequently hindered by hardware limitations. To address the issue of KV cache growth, various KV cache compression methods have been proposed. These methods can be broadly categorised into trainable approaches, which involve modifications to the model architecture Ainslie et al. (2023), or fine-tuning regime to inherently manage KV cache size Nawrot et al. (2024), and non-trainable approaches, which apply post-hoc compression techniques to reduce the cache footprint without altering the underlying model Li et al. (2024); Zhang et al. (2024); Ge et al. (2023a). While these methods have shown promise, they often involve complex algorithms or significant computational overhead, limiting their practicality.
We show that the $L_{2}$ norm of cached keys has a high correlation with attention scores. More specifically, we observe that a low $L_{2}$ norm of a key embedding usually leads to a high attention score during decoding. Based on this observation, we propose a simple and highly effective strategy for KV cache compression: keeping in memory only the keys with lowest $L_{2}$ norm, and the corresponding values. Unlike many existing methods, our heuristic can be applied off-the-shelf to any transformer-based decoder-only LLM without the need for additional training or significant modifications. More importantly, our method estimates the influence of cached key-value pairs without the need to compute the attention scores.
Our experimental results demonstrate that this heuristic allows maintaining model performance in language modelling tasks and in tasks that require the model to store and retrieve the most critical information, such as passkey retrieval (Mohtashami and Jaggi, 2023) and needle-in-a-haystack tasks (Kamradt, 2023).
Figure 2: Five heads at layer 9 of Llama2-7b. Attention score (top) and $L_{2}$ norm (bottom) are highly correlated. We observe similar patterns across most layers and for a wide range of inputs. More examples provided in Appendix E
2 Patterns in the Attention Matrix
In the following, we examine the attention scores on the language modelling task. By analysing the key embeddings and the attention distribution, we observe that key embeddings with low $L_{2}$ norm are often associated with higher attention scores. In Fig. 2, we provide an example using Llama-2-7b (Touvron et al., 2023), where the columns represent different heads, the first row presents the attention distribution over the KV pairs, and the second row presents the $L_{2}$ norm of each key embedding. We observe that the tokens with high attention scores, such as "<s>" and ".", have significantly lower $L_{2}$ norm values than others. While Xiao et al. (2024) already observed peaked attention distributions for specific tokens, and Darcet et al. (2024) pointed out the influence of high $L_{2}$ norm tokens on attention maps, we are the first, to the best of our knowledge, to point out the correlation between the $L_{2}$ norm of the key embeddings and attention score. Based on our observation, we consider the following research question: can we compress the KV cache based on the $L_{2}$ norm of the key embeddings?
An intuitive way to estimate the influence of compressing the KV cache is by examining the attention scores that are dropped due to the compression. In the following, we formally define this influence.
Given a prompt consisting of $n$ tokens $(x_{1},x_{2},...,x_{n})$, the LLM first encodes them into a KV cache— this step is referred to as the pre-filling phase. Then, the model autoregressively generates the next token $x_{n+1}$. When performing KV cache compression, some key-value pairs may be dropped and thus cannot be attended to. We define the attention loss caused by the compression as the sum of the attention scores associated with the dropped KV pairs:
$\mathcal{L}^{m}_{l,h}=\sum_{p\in D_{l,h}}a_{l,h,p},$ (1)
where $a_{l,h,p}$ is the attention score of the $p$-th token in the layer $l$, head $h$. In Eq. 1, $D_{l,h}$ denotes the positions of $m$ pairs of dropped KV, $|D_{l,h}|=m$, which depends on the compression method. An ideal compression algorithm aims to drop the KV pairs with the lowest attention scores, which will have less impact on the output. However, such attention scores are unavailable for a compression algorithm since it needs $x_{n+1}$ to query the full KV cache in advance. Instead, we drop KV pairs with the highest $L_{2}$ norm in key embeddings and use attention loss caused by ideal compression as the reference:
$\mathcal{Y}_{l,h}^{m}=\mathcal{L}_{l,h}^{m}-\mathcal{L}_{l,h}^{m,ref},$ (2)
where $\mathcal{L}_{l,h}^{m,ref}$ is the reference attention loss, and $\mathcal{Y}_{l,h}^{m}$ is a non-negative value. A lower $\mathcal{Y}_{l,h}^{m}$ indicates a lower difference and thus a higher correlation between the attention score and the $L_{2}$ norm. To measure the overall difference between ideal attention score-based compression and $L_{2}$ norm-based compression, we sum up the $\mathcal{Y}_{l,h}^{m}$ over different numbers of compressed KV pairs:
$\mathcal{Y}_{l,h}=\sum_{m=1}^{n}\mathcal{Y}_{l,h}^{m}.$ (3)
We name the $\mathcal{Y}_{l,h}$ as ALr, which indicates the Attention Loss for a compression method using ideal attention loss as Reference.
Figure 3: Perplexity of Llama2-7B on Wikipedia-en. Pruning the KV cache by removing tokens with $L_{2}$ norm has no impact on perplexity. The vertical line represents the maximum context length.
In Fig. 1, we plot the $\mathcal{Y}$ across layers and heads. We observe that heads in the first two layers and some middle layers around the 12th layer have relatively high $\mathcal{Y}$ values. The heads in other layers have lower $\mathcal{Y}$ values, indicating a high correlation between $L_{2}$ norm and attention score. By leveraging this correlation, we can compress the KV cache based on the $L_{2}$ norm of key embeddings. Optionally, we can skip the compression at the layers with low correlation. We show ablation experiments skipping layers in Appendix B.
3 Experiments
We evaluate our method on language modelling and two long-context modelling tasks, i.e., needle-in-a-haystack and passkey retrieval. Based on the observation supported by Fig. 1, the heads in the first two layers usually have a low correlation between $L_{2}$ norm and attention score, so we do not perform compression on these layers as default. We conduct experiments to investigate the impact of compression on different layers in Appendix B.
3.1 Language Modelling Task
For language modelling, we let the KV cache grow until a specific pre-defined length and subsequently start to discard the tokens with the highest $L_{2}$ norm. We show in Fig. 3 that evicting even up to the 50% of KV Cache does not impact perplexity. Perplexity increases, as expected, once we exceed the pre-training context length. We show more results, including next token accuracy in Appendix B.
To further verify that keys with low $L_{2}$ norm capture significant information, we test other eviction strategies, i.e. keeping tokens with highest $L_{2}$ norm and keeping random tokens. It is clear from Fig. 4 that discarding tokens with low $L_{2}$ impairs performance, even more so than random discarding, thus highlighting the importance of these low $L_{2}$ norm keys.
Figure 4: Compression with different strategies. It is clear that discarding tokens with low $L_{2}$ norms is detrimental to perplexity and delivers performance lower than random.
3.2 Long-Context Modelling Tasks
We evaluate the compression capability on two long context-modelling tasks, needle-in-a-haystack (Kamradt, 2023) and passkey retrieval (Mohtashami and Jaggi, 2023). In both tasks, the model needs to identify and retrieve the important information from a long context to generate correct answers. Thus, these tasks test the compression method’s ability to keep important KV pairs and drop redundant ones.
In Fig. 5 and Fig. 6, we present the experimental results of Llama-2-7b-80k (Fu et al., 2024). We present more experiments on other models in Appendix C.
As shown in Fig. 5, the model can sustain its performance on the needle-in-a-haystack task while compressing 30% of the KV cache, and maintain 99% accuracy when compressing 50% of the KV cache. Additionally, the model can achieve 100% accuracy on the passkey retrieval task even when compressing 90% of the KV cache, as shown in Fig. 6.
Moreover, we compare other eviction strategies, like keeping KV pairs with low $L_{2}$ norm, keeping KV pairs with high $L_{2}$ norm, and keeping random KV pairs. In Fig. 5 and Fig. 6, we observe that the model cannot answer correctly when keeping only high $L_{2}$ norm KV pairs, obtaining near zero and zero accuracy on the needle-in-a-haystack and passkey retrieval tasks, respectively. When we randomly compress the KV cache, the performance decreases significantly faster than keeping low $L_{2}$ norm KV pairs. The above analysis indicates that KV pairs with low $L_{2}$ norm are critical to generating the correct answer and thus contain important information.
Figure 5: Overall accuracy of llama-2-7b-80k on the needle-in-a-haystack task. More results in Appendix C.
Figure 6: Overall accuracy of llama-2-7b-80k on the passkey retrieval task. More results in Appendix C.
4 Related Works
Recently, various long-context LLMs, such as Gemini-Pro-1.5 (Reid et al., 2024), Claude-3 (Anthropic, 2024), and GPT4 (Achiam et al., 2023), have shown the promising capability to process hundred thousands of tokens in the context. The increased number of input lengths results in a high decoding latency; thus, there has been a growing interest in speeding up the decoding with long contexts. Some works propose efficient memory management strategies to reduce the IO time overheads, e.g., PageAttention (Kwon et al., 2023), Infinite-LLM (Lin et al., 2024) and vAttention (Prabhu et al., 2024). Another line of research focuses on compressing the KV cache to improve efficiency. DMC (Nawrot et al., 2024) compresses KV cache by dynamically merging tokens while requiring expensive continual pre-training. For fine-tuning free compression strategy, H2O (Zhang et al., 2024) identifies important KV pairs by leveraging the attention scores from all queries, FastGen (Ge et al., 2023b) leverages the different attention patterns in different heads for compression, and SnapKV (Li et al., 2024) selects KV pairs based on attention scores from user’s query. Unlike these works, our method only utilises the $L_{2}$ norm of embedding for compression without leveraging the attention information, and to the best of our knowledge, we are the first to find that the influence of a KV pair can be determined by $L_{2}$ norm. Previous work (Darcet et al., 2024) finds the hidden states with high $L_{2}$ norm usually aggregate more important and global information. On the other hand, our findings indicate that a low $L_{2}$ norm of key embedding generally results in a high attention score.
5 Conclusion
In this paper, we introduced a simple yet highly effective strategy for KV cache compression in LLMs based on the $L_{2}$ norm of key embeddings. We show that there is a significant correlation between the $L_{2}$ norm of a key embedding and its attention score. Leveraging this observation, our proposed method compresses the KV cache by retaining only those keys with the lowest $L_{2}$ norm. Our experimental results on various tasks show that our compression strategy maintains the predictive accuracy of the model while significantly reducing the memory footprint. Our approach is straightforward and can be applied directly to any transformer-based, decoder-only LLM.
6 Limitations
While our research provides valuable insights, our experiments and analyses were conducted exclusively using the Llama2-7b, Llama2-7b-32k and Llama2-7b-80k models. In future work, we will test on Llama2-13b and Llama2-70b, as well as the Llama3 family, to verify that our findings generalise to other models and architectures. Additionally, while we show that the $L_{2}$ norm played a significant role in our experiments, we do not have a comprehensive theoretical explanation for why this is the case. Understanding the underlying reasons behind the importance of the $L_{2}$ norm would require further theoretical exploration and empirical validation.
Appendix A Further Analysis
We discuss further the correlation between $L_{2}$ norm and attention scores. In Fig. 7, we show the ALr from Eq. 1 for two heads from layer 10 and layer 0. We see that in layer 0, the difference is larger due to a lower correlation.
In Fig. 8, we display $L_{2}$ norm and attention correlation across heads and layers using the original Llama2-7b and the long context Llama2-7b-32k and Llama2-7b-80k. We can see that patterns are quite consistent across all the models.
(a) Layer-7 Head-10, which has a high correlation between attention score and $L_{2}$-Norm.
(b) Layer-0 Head-0, which has a low correlation between attention score and $L_{2}$-Norm.
Figure 7: Attention loss of ideal compression and $L_{2}$ norm-based compression in Llama-2-7b-80k. The $x$-axis represents the compression ratio; the $y$-axis represents the attention loss (defined by Eq. 1) The results average over 1024 chunks on Wikipedia, with a length of 1024.
(a) Llama-2-7b.
(b) Llama-2-7b-80k
(c) Llama-2-7b-longlora-32k-ft
Figure 8: $\mathcal{Y}$ (defined by Eq. 3) of different models. The light square means a high difference between $L_{2}$-Norm. We observe that llama-2 and its fine-tuned long-context versions have a consistent pattern – the heads in the first two layers and middle layers are more different than others.
Appendix B More results on Language modelling task
In the following, we show results when performing compression only on layers that show a lower correlation between $L_{2}$ norm and attention score. We show in Fig. 12 that for language modelling tasks, the different layer drop has little impact on final accuracy and perplexity. The difference becomes significant only when the KV cache is pruned to retain only one thousand pairs. All experiments are averaged over 50 chunks from English Wikipedia.
Figure 9: Results on language modelling task when skipping the first layer.
Figure 10: Results on language modelling task when skipping the first two layers.
Figure 11: Results on language modelling task when skipping layers 0,1 and 12.
Figure 12: Skipping compression at different layers
Appendix C More Results on Long-Context Modelling Tasks
In addition to llama-2-7b-80k (Fu et al., 2024), we test the compression method using llama-2-7b-longlora-32k-ft (Chen et al., 2023) on the needle-in-a-haystack and passkey retrieval tasks. As shown in Fig. 13(a), we can see that compressing 30% of KV cache only results in a slight performance degradation on the needle-in-a-haystack task. We also observe that the performance even increases slightly when we compress 10% of KV cache. In figure Fig. 13(b), we observe that the llama-2-7b-longlora-32k-ft maintains 100% performance when compressing 80% of KV cache and only as a slight decrease when compressing 90% of KV cache. Furthermore, the model fails to generate correct answers if we compress KV pairs with low $L_{2}$ norm and keep high $L_{2}$ norm ones. The evaluation results of llama-2-7b-longlora-32k-ft are consistent with the llama-2-7b-80k, which further indicates the effectiveness of compressing KV cache using $L_{2}$ norm.
C.1 Analysis of Skipped Layers
As shown in Fig. 1, we find heads in the first two layers and the middle layers have a relatively low correlation between attention scores and $L_{2}$ norm. Thus, we conduct experiments to analyse the impact of skipping layers that have a low correlation for compression. As shown in Fig. 14(a) and Fig. 14(c), we observe that only skipping the first layer (layer-0) decreases the performance on the needle-in-a-haystack task significantly. We can see that skipping the first two layers (layer-0,1) has a similar performance compared to skipping the first three layers (layer-0,1,2). Furthermore, as shown in Fig. 14(b) and Fig. 14(d), only skipping the first layer can result in significant performance degradation. We also find that the compression ratio is not proportional to the overall accuracy of models in the passkey retrieval task when we compress the first layer, where the accuracy shows a U-shape curve regarding the compression ratio.
(a) Overall accuracy of Llama-2-7b-longlora-32k-ft on the needle-in-a-haystack task.
(b) Overall accuracy of Llama-2-7b-longlora-32k-ft on the passkey retrieval task.
Figure 13: Evaluation results of Llama-2-7b-longlora-32k-ft on the needle-in-a-haystack and passkey retrieval tasks.
(a) Overall accuracy of Llama-2-7b-80k on the needle-in-a-haystack task.
(b) Overall accuracy of Llama-2-7b-80k on the passkey retrieval task.
(c) Overall accuracy of Llama-2-7b-longlora-32k-ft on the needle-in-a-haystack task.
(d) Overall accuracy of Llama-2-7b-longlora-32k-ft on the passkey retrieval task.
Figure 14: Analysing of skipping different layers for compression.
(a) Llama-2-7b-80k, skip layer-0, compression ratio 30%
(b) Llama-2-7b-80k, skip layer-0 and layer-1, compression ratio 30%
(c) Llama-2-7b-80k, skip layer-0 and layer-1, compression ratio 20%
Figure 15: Detailed results of Llama-2-7b-80k on the needle-in-a-haystack task.
(a) Llama-2-7b-longlora-32k-ft, without compression
(b) Llama-2-7b-longlora-32k-ft, skip layer-0, compression ratio 30%
(c) Llama-2-7b-longlora-32k-ft, skip layer-0 and layer-1, compression ratio 30%
Figure 16: Detailed results of Llama-2-7b-longlora-32k-ft on the needle-in-a-haystack task.
(a) Llama-2-7b-80k, skip layer-0, compression ratio 90%
(b) Llama-2-7b-80k, skip layer-0 and layer-1, compression ratio 90%
Figure 17: Accuracy on the passkey retrieval. The $x$-axis presents the position of the passkey, and the $y$-axis presents the accuracy.
(a) Llama-2-7b-longlora-32k-ft, skip layer-0, compression ratio 90%
(b) Llama-2-7b-longlora-32k-ft, skip layer-0 and layer-1, compression ratio 90%
Figure 18: Accuracy on the passkey retrieval. The $x$-axis presents the position of the passkey, and the $y$-axis presents the accuracy.
Appendix D Experimental setup
In all experiments, we used the HuggingFace library and did not change the model’s default hyperparameters. For language modelling, results are averaged across 50 samples. The Fig. 7 and Fig. 1 are the average results of $1024$ examples with a chunk size of $1024$ using Wikipedia.
Appendix E More Visualizations
Figure 19: Attention maps in Llama2-7B
Figure 20: Norms of KV cache tokens in Llama2-7B
Figure 21: Attention maps in Llama2-7B
Figure 22: Norms of KV cache tokens in Llama2-7B
Figure 23: Attention maps in Llama2-7B
Figure 24: Norms of KV cache tokens in Llama2-7B